HUBBLE SPACE TELESCOPE DAILY REPORT #1588 PERIOD COVERED: 0000Z (UTC) 03/11/96 - 0000Z (UTC) 03/12/96 Daily Status Report as of 072/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. = During the second iteration of this proposal, at 071/0210Z, there was a = guide star acquisition failure, and the take data flag was down for the = entire period. HSTAR 5532 was written. Otherwise, the other = observations were completed as scheduled, and no problems were reported. 1.2 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy IR11095-02. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.3 Completed WF/PC-2 5990 (Globular Clusters in the Fornax Galaxies) The WF/PC-2 was used to search for globular clusters in galaxy = NGC-1379 which is a member of the Fornax Galactic Cluster. Because the = galaxies in the Fornax Cluster have different morphologies than those of = the Local Group, the globular cluster systems are also expected to show = differences. The first part of this proposal was affected by a guide = star acquisition failure at 071/0210Z when the search limit flag was = exceeded. A subsequent FHST map showed bad attitude, and an ARU/PRT was = performed at 071/0232Z. The reacquisition at 071/0346Z was successful. = HSATR 5532 describes these events. Otherwise, the observations were = executed as scheduled, and no further anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 5829 (The Activity of Periodic Comet = Schwassmann-Wachmann 1) The WF/PC-2 was used to make observations of periodic comet SW1 = (aka, Schwassmann-Wachmann 1) to try and detect CO fluorescence. The = observations were completed as planned, and no anomalies were noted. 1.5 Completed WF/PC-2 5913 (Low-Mass Binaries and Luminosity Fuction) The WF/PC-2 was used to make observations of the low-mass binary = system NGP009 (Late Type Class M dwarf stars). The intent is to = determine why the local stellar luminosity function is flat below = Mv=3D12, while that found from more distant stars falls by an order of = magnitude by Mv=3D14. The observations were executed as scheduled, and = no anomalies were noted. 1.6 Completed Two Sets of FOS/BL and HRS/2 5169 (High Resolution = Spectra of the Nucleus of NGC-4151: Cycle 4 High) The Faint Object Spectrograph (Blue Side) and the High Resolution = Spectrograph (Side Two) were used to make observations of the nucleus of = galaxies NGC-4151 and NGC-4151 to study the C IV, C II, and Si II lines. = During the first iteration, which used the FOS/BL, the Y-Speed test = failed at 071/110418Z, indicating that the FOS microprocessor load may = have been too heavy to perform all background tasks. HSTAR 5533 = documented this event. Then, during the second iteration, which utilized = the HRS/2 instrument, a one minute loss of lock occurred, beginning at = 071/144458Z. HSTAR 5534 was written. Otherwise, the observations were = executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 A loss of lock, of approximately a one minute duration, was = experienced at 071/144458Z. HSTAR 5534 was written. Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Operations Notes: A new FGS star selector angle bias calculation was installed = onboard the HST at approximately 071/1519Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH